OpenStudy (anonymous):

For all s<0, what happens to the value of (s^3-s^2)/(s^2)+s, as "s" decreases? A It decreases in proportion to s^3. B. It decreases in proportion to s. C. It remains constant. D. It increases in proportion to s^3. E. It increases in proportion to s.

OpenStudy (anonymous):

\[\frac{ s^3 +s^2 }{ s^2 + s }= \frac{ s^2 (s+1)}{s(s+1)}=s\] by factorising the top and bottom and then cancelling. So it will decrease in proportion to s.

OpenStudy (anonymous):

(It is in fact equal to s)

OpenStudy (anonymous):

thanx